OpenStudy (anonymous):

Use the method of substitution to solve the system of linear equations 8x+4y=8/5 y=5x-1

OpenStudy (anonymous):

plug in the value of Y into the equation, 8x+4y=8/5 instead of 4y you would put in the value of Y like this: 8x+4(5x-1)=8/5 then just solve(: do you get that or you need more help?

OpenStudy (anonymous):

Substitution means you can plug in one value of y for another. they made it easy for you by singling out the y, by y=5x-1 Plug y=5x-1 into the other equation, 8x+4(5x-1)=8/5. From there you can single out the value of x, and plug x back in to get y.

OpenStudy (anonymous):

okay what i have so far is 28x=9/5 is that right so far?

OpenStudy (anonymous):

I don't understand how you got 28x=9/5 You should make sure to factor out correctly. 4(5x-1)= 20x-4. 20x-4+8x= 28x-4=8/5. 28x= 8/5 +4. From there it should be easy to get the x. After getting x, plug into the original equations(hint, pick the easier one) and voila you got y.

OpenStudy (anonymous):

does x =1?

OpenStudy (anonymous):

oops i messed that up i think

OpenStudy (anonymous):

okay so 28x=8/5+4. so do i combine 8/5 and 4?